# Theoretical study of orbital torque: Dependence on ferromagnet species and nonmagnetic layer thickness

### v1

@@ Abstract
The manipulation of magnetization in ferromagnetic metals (FMs) through orbital torque (OT) has emerged as a promising route for energy-efficient magnetic devices without relying on heavy metals. While Ti and Cu are among the most extensively studied light nonmagnetic metals (NMs) for OT devices, theoretical calculations of the resulting torque have remained limited. Here, we present a systematic and quantitative theoretical study of current-induced torques in Ti/FM and Cu/FM (FM = Co, Ni) bilayers using semi-realistic tight-binding models derived from ab initio electronic structures. We find that the torque in Ti/FM is larger for Ni than for Co, but this trend does not necessarily hold in Cu/FM, revealing that the FM dependence of OT is not universal but varies with the orbital current source. Moreover, the dependence of OT on NM thickness clearly indicates its NM bulk origin in both Ti- and Cu-based systems. Notwithstanding, the quantitative characteristics of OT cannot be explained by a simplified picture based on the individual bulk properties of the NM or FM layers. These results provide microscopic insight and practical guidance for designing light-metal-based orbitronic devices.
@@ I Introduction
The manipulation of magnetization by electrical means has become a central topic in spintronics. Spin-orbit torques (SOTs)Manchon et al. (2019); Han et al. (2021); Shao et al. (2021) exploit efficient charge-to-spin conversion, typically achieved by the spin Hall effect (SHE) or the spin Rashba-Edelstein effect through the spin-orbit coupling (SOC) of a nonmagnetic metal (NM). The injected spin angular momentum exerts a torque on the magnetization of an adjacent ferromagnetic metal (FM) layer. Recently, an alternative mechanism, known as the orbital torque (OT),Go and Lee (2020); Ding et al. (2020); Kim et al. (2021); Lee et al. (2021a, b); Hayashi et al. (2023) has emerged as a new route to control magnetization through orbital angular momentum (OAM) transfer from a NM layer to a FM layer. OTs arise from an OAM current (or simply, orbital current) generated by the orbital Hall effect (OHE), or from an interfacial OAM accumulation by the orbital Rashba-Edelstein effect, both of which can occur even in the absence of SOC. Go et al. (2018); Salemi et al. (2019) For the transferred OAM to exert a torque on the magnetization of the FM layer, SOC must eventually intervene, since there is no direct exchange coupling between OAM and the local spin moment. Go and Lee (2020) Thus, in contrast to the SOT mechanism, in which the spin current source itself requires SOC to generate spin currents, the OT mechanism assigns the role of SOC in orbital-to-spin conversion to the layer that receives the orbital current, rather than to its source. This mechanism opens a distinct pathway to efficient torque generation using light NMs with weak SOC, drawing increasing attention to the emerging field of orbitronics. Go et al. (2021); Jo et al. (2024); Ando (2025)
@@
While a variety of light NMs have been examined for OT devices,Ando (2025) Ti Hayashi et al. (2023); Choi et al. (2023); Zhao et al. (2025); Xu et al. (2025a); Shin et al. (2025); Hayashi et al. (2025); Xu et al. (2025b); Zheng et al. (2025) and Cu An et al. (2016); Kim et al. (2021); Ding et al. (2020); Tazaki et al. (2020); Kim et al. (2023); Krishnia et al. (2024); Ding et al. (2024); Damerio and Avci (2025); Li and Pai (2025); Yi et al. (2025) are among the most extensively studied materials. Ti is a representative orbital Hall material, as experimentally confirmed through magneto-optical detection Choi et al. (2023) and transmission electron microscopy. Idrobo et al. (2024) The substantial orbital current generated by the OHE in Ti serves as a source of the OT in Ti/FM bilayers. Although the OHE itself does not require SOC, the OT mechanism relies on orbital-to-spin conversion mediated by SOC. To enhance this conversion efficiency, one typically employs a magnetic layer with stronger SOC Xu et al. (2025b) or inserts a thin heavy metal layer between the Ti and FM layers.Xu et al. (2025a); Shin et al. (2025); Zheng et al. (2025) In this regard, a pronounced dependence on the FM species is a hallmark feature of OT. Notably, even among the 3dd FMs, Ni has been reported to generate larger OTs than other 3dd FMs, which is attributed to its slightly more efficient orbital-to-spin conversion. Go et al. (2020); Lee et al. (2021a, b); Dutta and Tulapurkar (2022); Bose et al. (2023); Fukunaga et al. (2023); Moriya et al. (2024) This superiority of Ni has also been observed in Ti/FM bilayers. Hayashi et al. (2023, 2025) Another remarkable feature of Ti-based OT devices is the long characteristic length in the Ti-thickness dependence, reaching several tens of nanometers. Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025a); Hayashi et al. (2025) Although such long-range behavior is not yet fully understood, it indicates that the bulk Ti layer serves as the primary source of the OT.
@@
Addressing this question calls for a comparative theoretical study of OTs in NM/FM bilayers with various combinations of NMs and FMs. Given the long-range behavior of OTs, Go et al. (2023); Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025a); Hayashi et al. (2025); Bose et al. (2023); Fukunaga et al. (2023); Moriya et al. (2024) it is desirable to investigate sufficiently large bilayer systems, which poses a challenge for fully ab initio methods. Although several theoretical studies Go et al. (2020); Salemi, Berritta, and Oppeneer (2021); Go et al. (2023); Nikolaev et al. (2024); Devda et al. (2025); Pezo et al. (2025) have examined the OT in various heterostructures, a systematic theoretical investigation of the FM dependence of OT in light NM/FM bilayers has, to our knowledge, not yet been performed. It is worth noting that the FM dependence could also originate from a self-induced SOT within the FM layer, also known as the anomalous SOT. Wang et al. (2019) Examining the dependence of the torque on material combination and layer thickness will provide valuable insight into distinguishing OT from other possible contributions.
In this work, we perform a systematic theoretical study of current-induced torques in NM/FM bilayers, where NM = Ti or Cu and FM = Co or Ni. By employing semi-realistic tight-binding models with parameters extracted from ab initio calculations, we achieve relatively large system sizes—up to 50 atomic layers ($\gtrsim 10$ nm). For Ti/FM bilayers, we find that the torque is larger for Ni, consistent with experimental observations. Hayashi et al. (2023, 2025) For Cu-based systems, we focus on pristine Cu/FM bilayers without oxidation. Although OT in these systems is generally assumed to be negligible, our results reveal non-negligible OT signals, indicating a finite OHE in Cu as predicted theoretically Jo, Go, and Lee (2018); Salemi and Oppeneer (2022); Go et al. (2024) and observed experimentally. Rothschild et al. (2022); Ben Tal et al. (2025) Interestingly, the trend of FM dependence reverses in Cu/FM systems under certain conditions, with Co producing a stronger torque. This finding demonstrates that the FM dependence of OT is not universal but can vary depending on the orbital current source, even when only intrinsic effects are considered. We also examine the NM thickness dependence, confirming the bulk NM origin of OTs in both Ti/FM and Cu/FM systems. In Cu/Ni, the bulk Cu contribution exhibits the opposite sign to that in Cu/Co, cautioning that OT cannot be directly predicted from the bulk orbital Hall conductivity of the NM. Consequently, our results demonstrate that the characteristics of OT cannot be explained solely by the individual bulk properties of the NM or FM layers, but depend critically on their material combination. These findings provide microscopic insight into the design of efficient orbitronic devices.
@@ II.1 Tight-binding models based on ab initio electronic structures
We construct semi-realistic tight-binding models for NM/FM systems using the Wannier functions extracted from ab initio band structures. The bulk band structures of the NMs (Ti and Cu) and FMs (Co and Ni) are obtained using the full-potential linearized augmented plane-wave code FLEURfle ; Wortmann et al. (2023) The Perdew-Burke-Ernzerhof functional Perdew, Burke, and Ernzerhof (1996) within the generalized gradient approximation is used for the exchange-correlation functional. All systems are assumed to have the face-centered cubic (fcc) structure, and the lattice constants of the FMs are set equal to those of the corresponding NMs. Specifically, for Co and Ni, two sets of band structures are independently calculated using either the lattice constant of Ti Chakraborty et al. (2011) ($a_{\mathrm{Ti}}=4.16$ Å) or that of Cu Straumanis and Yu (1969) ($a_{\mathrm{Cu}}=3.61$ Å), assuming that the FM layer adopts the same structure as the adjoining NM layer in each NM/FM bilayer. The magnetization of Ni and Co is aligned along the normal vector to the fcc(111) plane. SOC is treated within the second variation scheme, Li et al. (1990) and the Brillouin zone is sampled using a $12\times 12\times 12$ Monkhorst-Pack $\mathbf{k}$-mesh. Monkhorst and Pack (1976)
@@
II.2 Linear-response calculation of the current-induced torque
@@
where $t_{ij}$ is the component of the linear-response tensor $\mathbf{t}$, referred to as the torkance, and $i$ and $j$ denote Cartesian components. In the Kubo formalism, $t_{ij}$ is decomposed into intraband and interband contributions, $t_{ij}=t_{ij}^{\mathrm{intra}}+t_{ij}^{\mathrm{inter}}$, each of which is evaluated as follows: Freimuth, Blügel, and Mokrousov (2014); Go et al. (2020)
@@
$\times\langle\psi_{n\mathbf{k}}|\hat{T}_{i}|\psi_{n\mathbf{k}}\rangle\langle\psi_{n\mathbf{k}}|\hat{v}_{j}|\psi_{n\mathbf{k}}\rangle,$
@@
$\times\mathrm{Im}\left[\frac{\langle\psi_{n\mathbf{k}}|\hat{T}_{i}|\psi_{m\mathbf{k}}\rangle\langle\psi_{m\mathbf{k}}|\hat{v}_{j}|\psi_{n\mathbf{k}}\rangle}{(\epsilon_{n\mathbf{k}}-\epsilon_{m\mathbf{k}})(\epsilon_{n\mathbf{k}}-\epsilon_{m\mathbf{k}}+i\Gamma)}\right],$
where $e>0$ is the elementary charge, $f_{n\mathbf{k}}=1/[e^{(\epsilon_{n\mathbf{k}}-\mu)/k_{\mathrm{B}}T}+1]$ is the Fermi-Dirac distribution function, with the chemical potential $\mu$ and $k_{\mathrm{B}}T=25$ meV, $\hat{\mathbf{v}}=\frac{i}{\hbar}[\hat{H},\hat{\mathbf{r}}$] is the velocity operator, and $\hat{\mathbf{T}}=\frac{i}{\hbar}[\hat{H}_{\mathrm{XC}},\hat{\mathbf{S}}$] is the exchange torque operator. The effect of disorder is approximately described by a constant energy broadening $\Gamma$. The OAM current or accumulation is not explicitly considered in the torque calculations using Eqs. (4) and (5), but its contribution to the OT is implicitly included. Go et al. (2020) The intraband term $t_{ij}^{\mathrm{intra}}$ in Eq. (4) corresponds to the (Drude-like) Boltzmann contribution, which scales with the momentum relaxation time $\tau\propto 1/\Gamma$. In contrast, the interband term $t_{ij}^{\mathrm{inter}}$ in Eq. (5) arises from the Berry curvature mechanism and represents an intrinsic contribution. Notably, $t_{ij}^{\mathrm{intra}}$ and $t_{ij}^{\mathrm{inter}}$ are odd and even under time-reversal, respectively. Symmetry considerations indicate that $t_{ij}^{\mathrm{intra}}$ gives rise to the field-like torque, whereas $t_{ij}^{\mathrm{inter}}$ leads to the damping-like torque. Go et al. (2020) Therefore, we calculate only $t_{yx}^{\mathrm{inter}}\equiv t_{yx}$ to describe the damping-like torque induced by an electric field applied along $\hat{\mathbf{x}}$.

### v2

@@ Abstract
The manipulation of magnetization in ferromagnetic metals (FMs) through orbital torque (OT) has emerged as a promising route for energy-efficient magnetic devices without relying on heavy metals. While Ti and Cu are among the most extensively studied light nonmagnetic metals (NMs) for OT devices, theoretical calculations of the resulting torque have remained limited. Here, we present a systematic and quantitative theoretical study of current-induced torques in Ti/FM and Cu/FM (FM = Co, Ni) bilayers using realistic tight-binding models derived from ab initio electronic structures. We find that the torque in Ti/FM is larger for Ni than for Co, but this trend does not necessarily hold in Cu/FM, revealing that the FM dependence of OT is not universal but varies with the orbital current source. Moreover, the dependence of OT on NM thickness clearly indicates its NM bulk origin in both Ti- and Cu-based systems. Notwithstanding, the quantitative characteristics of OT cannot be explained by a simplified picture based on the individual bulk properties of the NM or FM layers. These results provide microscopic insight and practical guidance for designing light-metal-based orbitronic devices.
@@ I Introduction
The manipulation of magnetization by electrical means has become a central topic in spintronics. Spin-orbit torques (SOTs)Manchon et al. (2019); Han et al. (2021); Shao et al. (2021) exploit efficient charge-to-spin conversion, typically achieved by the spin Hall effect (SHE) or the spin Rashba-Edelstein effect through the spin-orbit coupling (SOC) of a nonmagnetic metal (NM). The injected spin angular momentum exerts a torque on the magnetization of an adjacent ferromagnetic metal (FM) layer. Recently, an alternative mechanism, known as the orbital torque (OT),Go and Lee (2020); Ding et al. (2020); Kim et al. (2021); Lee et al. (2021a, b); Hayashi et al. (2023) has emerged as a new route to control magnetization through orbital angular momentum (OAM) transfer from a NM layer to a FM layer. OTs arise from an OAM current (or simply, orbital current) generated by the orbital Hall effect (OHE), or from an interfacial OAM accumulation by the orbital Rashba-Edelstein effect, both of which can occur even in the absence of SOC. Go et al. (2018); Salemi et al. (2019) For the transferred OAM to exert a torque on the magnetization of the FM layer, SOC must eventually intervene, since there is no direct exchange coupling between OAM and the local spin moment. Go and Lee (2020) Thus, in contrast to the SOT mechanism, in which the spin current source itself requires SOC to generate spin currents, the OT mechanism assigns the role of SOC in orbital-to-spin conversion to the layer that receives the orbital current, rather than to its source. This mechanism opens a distinct pathway to efficient torque generation using light NMs with weak SOC, drawing increasing attention to the emerging field of orbitronics. Go et al. (2021); Jo et al. (2024); ,Kazuya (2025)
@@
While a variety of light NMs have been examined for OT devices,,Kazuya (2025) Ti Hayashi et al. (2023); Choi et al. (2023); Zhao et al. (2025); Xu et al. (2025b); Shin et al. (2025); Hayashi et al. (2025); Xu et al. (2025a); Zheng et al. (2025) and Cu An et al. (2016); Kim et al. (2021); Ding et al. (2020); Tazaki et al. (2020); Kim et al. (2023); Krishnia et al. (2024); Ding et al. (2024); Damerio and Avci (2025); Li and Pai (2025); Yi et al. (2025) are among the most extensively studied materials. Ti is a representative orbital Hall material, as experimentally confirmed through magneto-optical detection Choi et al. (2023) and transmission electron microscopy. Idrobo et al. (2024) The substantial orbital current generated by the OHE in Ti serves as a source of the OT in Ti/FM bilayers. Although the OHE itself does not require SOC, the OT mechanism relies on orbital-to-spin conversion mediated by SOC. To enhance this conversion efficiency, one typically employs a magnetic layer with stronger SOC Xu et al. (2025a) or inserts a thin heavy metal layer between the Ti and FM layers.Xu et al. (2025b); Shin et al. (2025); Zheng et al. (2025) In this regard, a pronounced dependence on the FM species is a hallmark feature of OT. Notably, even among the 3dd FMs, Ni has been reported to generate larger OTs than other 3dd FMs, which is attributed to its slightly more efficient orbital-to-spin conversion. Go et al. (2020); Lee et al. (2021a, b); Dutta and Tulapurkar (2022); Bose et al. (2023); Fukunaga et al. (2023); Moriya et al. (2024) This superiority of Ni has also been observed in Ti/FM bilayers. Hayashi et al. (2023, 2025) Another remarkable feature of Ti-based OT devices is the long characteristic length in the Ti-thickness dependence, reaching several tens of nanometers. Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025b); Hayashi et al. (2025) Although such long-range behavior is not yet fully understood, it indicates that the bulk Ti layer serves as the primary source of the OT.
@@
Addressing this question calls for a comparative theoretical study of OTs in NM/FM bilayers with various combinations of NMs and FMs. Given the long-range behavior of OTs, Go et al. (2023); Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025b); Hayashi et al. (2025); Bose et al. (2023); Fukunaga et al. (2023); Moriya et al. (2024) it is desirable to investigate sufficiently large bilayer systems, which poses a challenge for fully ab initio methods. Although several theoretical studies Go et al. (2020); Salemi et al. (2021); Go et al. (2023); Nikolaev et al. (2024); Devda et al. (2025); Pezo et al. (2025) have examined the OT in various heterostructures, a systematic theoretical investigation of the FM dependence of OT in light NM/FM bilayers has, to our knowledge, not yet been performed. It is worth noting that the FM dependence could also originate from a self-induced SOT within the FM layer, also known as the anomalous SOT. Wang et al. (2019) Examining the dependence of the torque on material combination and layer thickness will provide valuable insight into distinguishing OT from other possible contributions.
In this work, we perform a systematic theoretical study of current-induced torques in NM/FM bilayers, where NM = Ti or Cu and FM = Co or Ni. By employing realistic tight-binding models with parameters extracted from ab initio calculations, we achieve relatively large system sizes—up to 50 atomic layers ($\gtrsim 10$ nm). For Ti/FM bilayers, we find that the torque is larger for Ni, consistent with experimental observations. Hayashi et al. (2023, 2025) For Cu-based systems, we focus on pristine Cu/FM bilayers without oxidation. Although OT in these systems is generally assumed to be negligible, our results reveal non-negligible OT signals, indicating a finite OHE in Cu as predicted theoretically Jo et al. (2018); Salemi and Oppeneer (2022); Go et al. (2024) and observed experimentally. Rothschild et al. (2022); Ben Tal et al. (2025) Interestingly, the trend of FM dependence reverses in Cu/FM systems under certain conditions, with Co producing a stronger torque. This finding demonstrates that the FM dependence of OT is not universal but can vary depending on the orbital current source, even when only intrinsic effects are considered. We also examine the NM thickness dependence, confirming the bulk NM origin of OTs in both Ti/FM and Cu/FM systems. In Cu/Ni, the bulk Cu contribution exhibits the opposite sign to that in Cu/Co, cautioning that OT cannot be directly predicted from the bulk orbital Hall conductivity of the NM. Consequently, our results demonstrate that the characteristics of OT cannot be explained solely by the individual bulk properties of the NM or FM layers, but depend critically on their material combination. These findings provide microscopic insight into the design of efficient orbitronic devices.
@@ II.1 Tight-binding models based on ab initio electronic structures
We construct realistic tight-binding models for NM/FM systems using the Wannier functions extracted from ab initio band structures. The bulk band structures of the NMs (Ti and Cu) and FMs (Co and Ni) are obtained using the full-potential linearized augmented plane-wave code FLEUR70; D. Wortmann, G. Michalicek, N. Baadji, M. Betzinger, G. Bihlmayer, J. Bröder, T. Burnus, J. Enkovaara, F. Freimuth, C. Friedrich, C. Gerhorst, S. Granberg Cauchi, U. Grytsiuk, A. Hanke, J. Hanke, M. Heide, S. Heinze, R. Hilgers, H. Janssen, D. A. Klüppelberg, R. Kovacik, P. Kurz, M. Lezaic, G. K. H. Madsen, Y. Mokrousov, A. Neukirchen, M. Redies, S. Rost, M. Schlipf, A. Schindlmayr, M. Winkelmann, and S. Blügel (2023) The Perdew-Burke-Ernzerhof functional Perdew et al. (1996) within the generalized gradient approximation is used for the exchange-correlation functional. All systems are assumed to have the face-centered cubic (fcc) structure, and the lattice constants of the FMs are set equal to those of the corresponding NMs. Specifically, for Co and Ni, two sets of band structures are independently calculated using either the lattice constant of Ti Chakraborty et al. (2011) ($a_{\mathrm{Ti}}=4.16$ Å) or that of Cu Straumanis and Yu (1969) ($a_{\mathrm{Cu}}=3.61$ Å), assuming that the FM layer adopts the same structure as the adjoining NM layer in each NM/FM bilayer. The magnetization of Ni and Co is aligned along the normal vector to the fcc(111) plane. SOC is treated within the second variation scheme, Li et al. (1990) and the Brillouin zone is sampled using a $12\times 12\times 12$ Monkhorst-Pack $\mathbf{k}$-mesh. Monkhorst and Pack (1976)
@@
II.2 Theoretical description of the torque
To study OT, we employ the continuity equation for angular momentum transport in a general magnetic system described by the Hamiltonian $\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{XC}}+\hat{H}_{\mathrm{SO}}$. By combining the two equations of motion for spin and OAM, which are coupled to each other via SOC, one obtains the continuity equation for the total angular momentum operator $\hat{\mathbf{J}}=\hat{\mathbf{L}}+\hat{\mathbf{S}}$ as follows: Haney and Stiles (2010); Go et al. (2020)
$\frac{\partial\hat{J}_{i}(\mathbf{r})}{\partial t}=-\nabla\cdot\hat{\mathbf{Q}}^{J_{i}}(\mathbf{r})+\hat{T}_{i}^{\mathrm{XC}}(\mathbf{r})+\hat{T}_{i}^{\mathrm{CF}}(\mathbf{r}).$ (3)
[…]
$\hat{\mathbf{T}}^{\mathrm{XC}}$ $=\frac{i}{\hbar}[\hat{H}_{\mathrm{XC}},\hat{\mathbf{S}}$], (4)
$\hat{\mathbf{T}}^{\mathrm{CF}}$ $=\frac{i}{\hbar}[\hat{H}_{\mathrm{0}},\hat{\mathbf{L}}$] (5)
are the exchange torque and the crystal field torque operators, respectively. Note that $\hat{H}_{\mathrm{0}}$ in Eq. (5) contains the rotationally invariant kinetic energy ($\frac{-\hbar^{2}\nabla^{2}}{2m}$) and the crystal potential energy, but only the latter does not commute with $\hat{\mathbf{L}}$ and therefore contributes to $\hat{\mathbf{T}}^{\mathrm{CF}}$.
Equation (3) relates the local accumulation rate of angular momentum to the divergence (or flux) of the angular momentum current, e.g., generated by orbital and spin Hall effects. The torque terms act as sources or sinks that violate the conservation of total angular momentum in an electronic system, describing the transfer of angular momentum from conduction electrons to other degrees of freedom. The exchange torque $\hat{\mathbf{T}}^{\mathrm{XC}}$ describes transfer of angular momentum to the local magnetic moment via exchange coupling. The magnetization dynamics probed in torque experiments (including SOT, OT, etc.) is captured by this term. Therefore, we mainly focus on $\hat{\mathbf{T}}^{\mathrm{XC}}$ throughout this paper. The crystal field torque $\hat{\mathbf{T}}^{\mathrm{CF}}$ describes transfer of angular momentum to the lattice via the crystal field, i.e., quenching of OAM. Although this term is often disregarded, Eq. (3) explicitly shows that it plays an important role in angular momentum transport by connecting the non-conserved angular momentum current $\hat{\mathbf{Q}}^{\mathbf{J}}$ to experimentally relevant observables such as $\hat{\mathbf{J}}$ and $\hat{\mathbf{T}}^{\mathrm{XC}}$. Thus, the conventional definition of angular momentum current Tanaka et al. (2008); Kontani et al. (2009); Go et al. (2018); Jo et al. (2018); Salemi and Oppeneer (2022) alone is insufficient to determine the angular momentum accumulation or the torque. Because our approach directly computes $\hat{\mathbf{T}}^{\mathrm{XC}}$ exerted on the magnetization, this issue associated with defining the angular momentum current does not affect our analysis. In the presence of SOC, a complete separation of spin and orbital contributions to $\hat{\mathbf{T}}^{\mathrm{XC}}$ is generally nontrivial. Nonetheless, we will demonstrate that OT is the dominant mechanism in our systems.
II.3 Linear-response calculation of the current-induced torque
@@
where $t_{ij}$ is the component of the linear-response tensor $\mathbf{t}$, referred to as the torkance, and $i$ and $j$ denote Cartesian components. In the Kubo formalism, $t_{ij}$ is decomposed into intraband and interband contributions, $t_{ij}=t_{ij}^{\mathrm{intra}}+t_{ij}^{\mathrm{inter}}$, each of which is evaluated as follows: Freimuth et al. (2014); Go et al. (2020)
@@
$\times\langle\psi_{n\mathbf{k}}|\hat{T}_{i}^{\mathrm{XC}}|\psi_{n\mathbf{k}}\rangle\langle\psi_{n\mathbf{k}}|\hat{v}_{j}|\psi_{n\mathbf{k}}\rangle,$
@@
$\times\mathrm{Im}\left[\frac{\langle\psi_{n\mathbf{k}}|\hat{T}_{i}^{\mathrm{XC}}|\psi_{m\mathbf{k}}\rangle\langle\psi_{m\mathbf{k}}|\hat{v}_{j}|\psi_{n\mathbf{k}}\rangle}{(\epsilon_{n\mathbf{k}}-\epsilon_{m\mathbf{k}})(\epsilon_{n\mathbf{k}}-\epsilon_{m\mathbf{k}}+i\Gamma)}\right],$
where $e>0$ is the elementary charge, $f_{n\mathbf{k}}=1/[e^{(\epsilon_{n\mathbf{k}}-\mu)/k_{\mathrm{B}}T}+1]$ is the Fermi-Dirac distribution function, with the chemical potential $\mu$ and $k_{\mathrm{B}}T=25$ meV, $\hat{\mathbf{v}}=\frac{i}{\hbar}[\hat{H},\hat{\mathbf{r}}$] is the velocity operator, and $\hat{\mathbf{T}}^{\mathrm{XC}}$ is the exchange torque operator defined by Eq. (4). The effect of disorder is approximately described by a constant energy broadening $\Gamma$. The OAM current or accumulation is not explicitly considered in the torque calculations using Eqs. (7) and (8), but its contribution to the OT is implicitly included (see Sec. II.2). The intraband term $t_{ij}^{\mathrm{intra}}$ in Eq. (7) corresponds to the (Drude-like) Boltzmann contribution, which scales with the momentum relaxation time $\tau\propto 1/\Gamma$. In contrast, the interband term $t_{ij}^{\mathrm{inter}}$ in Eq. (8) arises from the Berry curvature mechanism and represents an intrinsic contribution. Notably, $t_{ij}^{\mathrm{intra}}$ and $t_{ij}^{\mathrm{inter}}$ are odd and even under time-reversal, respectively. Symmetry considerations indicate that $t_{ij}^{\mathrm{intra}}$ gives rise to the field-like torque, whereas $t_{ij}^{\mathrm{inter}}$ leads to the damping-like torque. Go et al. (2020) Therefore, we calculate only $t_{yx}^{\mathrm{inter}}\equiv t_{yx}$ to describe the damping-like torque induced by an electric field applied along $\hat{\mathbf{x}}$.